What is the sum of the solutions of the equation 3x^2+11x=4?
Answer by checkley71(8403)

3X^2+11X=4
3X^2+11X-4=0
(3X-1)(X+4)=0
3X-1=0
3X=1
X=1/3 SOLUTION
X+4=0
X=-4